OpenStudy (anonymous):

Which statement is true about the difference 2/7- /28? A. It is rational and equal to -2. B. It is rational and equal to 0. C. It is irrational and equal to -2 /7 D. It is irrational and equal to /7 THE SLASH SIGNS ARE SUPPOSED TO BE SQUARE ROOTS FYI :)

OpenStudy (anonymous):

What do you think?

OpenStudy (anonymous):

C? idk why tho

OpenStudy (anonymous):

Maybe this will help \( \huge 2\sqrt{7}- \sqrt{28} \) \( \huge \sqrt{4} \sqrt{7} = \sqrt{28}\) \( \huge 2\sqrt{7}- \sqrt{4} \sqrt{7} \) \( \huge 2\sqrt{7}- 2 \sqrt{7} = ? \)

OpenStudy (anonymous):

oohohhh lol B

OpenStudy (anonymous):

Good Job!!!!

OpenStudy (anonymous):

lol thnx so much :)

OpenStudy (anonymous):

YW